HUBBLE SPACE TELESCOPE DAILY REPORT #2200 PERIOD COVERED: 0000Z (UTC) 09/01/98 - 0000Z (UTC) 09/02/98 Daily Status Report as of 245/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of NIC/1/2/3 7899 (An IR Search For faint Companions to Pleiades Very Low-Mass Stars and Brown Dwarfs) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of CFHT-PL-15 and ROQUE15. This is part of a infrared campaign to search for faint companions to Pleiades very low-mass stars and brown dwarfs. The Pleiades has been the best site for brown dwarf {BD} searches, with half of the presently "confirmed" BDs being Pleiades members. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Ten Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.3 Completed Five Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.4 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed NIC/1/2/3 7827 (The Structure of Circumbinary Material in Multiple T-Tauri Systems) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the T-Tauri class DD-Tau to study the structure of its circumbinary material. Recent studies of T-Tauri populations in nearby star forming clouds have shown that the majority of stars exist in multiple {binary or higher} systems. The observations were completed as planned, and no problems were reported. 1.7 Completed FOC/96 6510 (The Spatially Extended Featureless Continuum Source of Cygnus A: Multicolor Imaging in Total Flux and Polarization) The Faint Object Camera (f/96) was used to make observations of the spatially extended featureless continuum source of the powerful prototype classical double radio galaxy Cygnus-A. Ground based data and several brief HST images show that the nuclear featureless continuum {FC} source is fully resolved spatially. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Nine Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.9 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.10 Completed NIC/2 7329 (The Nature of the Damped LyAlpha Absorbers -- A New Study of Young Galaxies) The Near Infrared Camera (Camera Two) was used to make observations of the young galaxy Q1610+2806 in order to study the nature of its damped Lyman-Alpha Absorbers (DLAs). The damped Lyman-Alpha absorption line systems, occasionally found in quasar spectra, provide one of only a few practical methods of detecting and studying significant numbers of galaxies at high redshifts. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: The SSR EDAC Error Counter was cleared one time (ROP SR-1). A correctable EDAC Type 1664 SSR Error was received (Area 13100-14800) at 245/0244Z. A SSRDIAG.SPT was performed and was sent to the DMS SE's public account. HSTAR 6644 was written. The NICMOS EDAC Single Bit Error Counter was incremented to eight at 245/0647Z. The NICMOS Internal log was dumped at 1040Z and the NSBITERR limit was adjusted to eight per Ops Note 13807. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH